OpenStudy (anonymous):

Factor y 2 + 5y + 6. A) (y + 1)(y + 6) B) (y + 2)(y + 3) C) (y + 2)(y + 4)

OpenStudy (anonymous):

thankyou

OpenStudy (aaronandyson):

y^2 + 5y + 6 (find two numbers when added gives 5 and when multipied gives )

OpenStudy (aaronandyson):

the number are 2 and 3

OpenStudy (aaronandyson):

so y^2 + 5y + 6 y^2 +2y + 3y + 6

OpenStudy (aaronandyson):

take the common factors

OpenStudy (aaronandyson):

so it would be (y+ 2)(y + 3) as y^2 + 2y take y common so it would be y + 2

OpenStudy (aaronandyson):

hope you understood

OpenStudy (anonymous):

yes i do thanks @AaronAndyson

OpenStudy (aaronandyson):

the link given by @Starr_DynastyT is wrong

OpenStudy (aaronandyson):

and np @Lakeekee_

OpenStudy (aaronandyson):

more questions?

OpenStudy (anonymous):

um yea i have like 2 more can you help @AaronAndyson

OpenStudy (aaronandyson):

sure

OpenStudy (anonymous):

The area of a rectangle is found by multiplying the length by the width: A = lw. A certain rectangle has an area of x 2 + 7x + 12. Factor the trinomial to find the length and width of the rectangle.

OpenStudy (aaronandyson):

is it x^2 + 7x + 12?

OpenStudy (anonymous):

yes

OpenStudy (aaronandyson):

x^2 + 7x + 12 = x^2 + 3x + 4x + 12 = x(x + 3) + 4(x + 3) = (x + 3)(x + 4)

OpenStudy (anonymous):

thankyou and I have one more

OpenStudy (aaronandyson):

ask :)

OpenStudy (anonymous):

these are the answer choices to the problem you just did: l = x + 6; w = x + 2 l = x + 5; w = x + 7 l = x + 4; w = x + 3 which one is the correct one

OpenStudy (aaronandyson):

the last one